### v4

# GORAG: Graph-based Online Retrieval Augmented Generation for Dynamic Few-shot Social Media Text Classification

Abstract.
Text classification is vital for Web for Good applications like hate speech and misinformation detection. However, traditional models (e.g., BERT) often fail in dynamic few-shot settings where labeled data are scarce, and target labels frequently evolve. While Large Language Models (LLMs) show promise in few-shot settings, their performance is often hindered by increased input size in dynamic evolving scenarios. To address these issues, we propose GORAG, a Graph-based Online Retrieval-Augmented Generation framework for dynamic few-shot text classification. GORAG constructs and maintains a weighted graph of keywords and text labels, representing their correlations as edges. To model these correlations, GORAG employs an edge weighting mechanism to prioritize the importance and reliability of extracted information and dynamically retrieves relevant context using a tailored minimum-cost spanning tree for each input. Empirical evaluations show GORAG outperforms existing approaches by providing more comprehensive and precise contextual information. Our code is released at: https://github.com/Wyb0627/GORAG.
Large Language Model, Retrieval-Augmented Generation, Few-shot Text Classification, Hate Speech Detection, Social Media
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3793011isbn: 979-8-4007-2307-0/2026/04ccs: Information systems Social networksccs: Information systems Clustering and classificationccs: Information systems Retrieval models and ranking
1. Introduction
Figure 1. An example of DFSTC task with 2 rounds. In Round 1, the text is classified as Threatening Language. In Round 2, another text is classified as a new label: Identity Insults.
Text content classification is a fundamental task for various social media applications, such as hate speech detection (Geissler et al., 2025; Hee et al., 2024; Chen et al., 2024b; Djuric et al., 2015; Subramanyam et al., 2024; Mandal et al., 2024; Cao et al., 2024; Ji et al., 2024; Guo et al., 2024), misinformation/disinformation detection (Luo et al., 2024; Nesterov et al., 2024; Liu et al., 2025; Toughrai et al., 2025; Cinus et al., 2025; Ashkinaze et al., 2024; Chen et al., 2022; Shang et al., 2024; Ye et al., 2024) , and suicide risk detection (Kang et al., 2024; Li et al., 2025; Naseem et al., 2025; Ge et al., 2025). In recent years, Neural Network (NN)-based models (Johnson and Zhang, 2015; Liu et al., 2016; Dieng et al., 2017; Wang et al., 2017; Johnson and Zhang, 2017; Peters and others, 2018; Wang et al., 2018; Yang et al., 2019), such BERT (Devlin, 2018) and RoBERTa (Liu, 2019), have demonstrated impressive performance on social media text content classification tasks. However, the effectiveness of these NN-based approaches (Johnson and Zhang, 2015; Liu et al., 2016; Dieng et al., 2017; Wang et al., 2017; Johnson and Zhang, 2017; Peters and others, 2018; Wang et al., 2018; Yang et al., 2019) depends on abundant labeled training data, which requires significant time and human efforts (Meng et al., 2018; Nguyen and Rudra, 2024). Furthermore, in real-world applications, such as Abuse Detection (Vidgen et al., 2021), new kinds of social media abuse can frequently appear with the changing needs of user analytics and progress in social research development, introducing the need to classify texts into new, extra labels (Xia et al., 2021), leading to the Dynamic Few-shot Social media Text Classification (DFSTC) task. As shown in Figure 1, the DFSTC task starts with an initial set of classes (e.g., Threatening Language). As new classes like Identity Insults emerge in later rounds, it becomes challenging to obtain extensive labeled data for these new categories. Models must therefore adapt to such changes and accurately classify new examples with limited labeled data. Developing methods to dynamically classify text with limited labeled data remains a crucial and open problem.
Depending on the technique for the DFSTC task, current models can be mainly categorized into two types, i.e., data augmentation-based models and RAG-based models. Firstly, data augmentation-based models (Meng et al., 2018, 2020; Xia et al., 2021) create additional data for fine-tuning small language models (e.g., RoBERTa (Liu, 2019)), by mixing the pairs of the few-shot labeled data and assigning mixed labels indicating the validity for these created data based on the labels of each data pairs (Xia et al., 2021). However, due to the limited labeled data, and the label names may not always be available; furthermore, the additional data created can have very limited patterns. Consequently, the classification models trained with these created data may over-fit and are not generalizable (Liu et al., 2023, 2024a; Lei et al., 2023).
Recently, Large Language Models (LLMs) (Sun et al., 2021; Jiang et al., 2023; Touvron et al., 2023; Achiam et al., 2023; Yang et al., 2024a) have excelled in social media text classification through their superior understanding abilities. However, they often struggle with task-specific knowledge (Gao et al., 2023; Zhang et al., 2023) and lack accurate reasoning clues to explain responses (Peng et al., 2024; Wang et al., 2024b; Dong et al., 2025; Qian et al., 2025). In domains like hate speech or misinformation detection, these flaws can lead to undetected harmful content or wrongful penalties. Consequently, Long Context RAG models (Chen et al., 2023; Sarthi et al., 2024; Xia et al., 2021; Yoon et al., 2024) were proposed to integrate retrieved side information (e.g., label descriptions) as external clues for more accurate classification. However, the incorporation of side information can further increase the input size (Chen et al., 2023), which often overwhelms critical details with noises. As the key clues of hate speech or misinformation are usually short sentences within long contexts, the ignorance of fine-grained, entity-level information can degrade LLM classification performance. Although proprietary LLMs (Hurst et al., 2024) can handle such extended inputs, they inevitably suffer from high token costs.
To achieve finer-grained RAG, and better preserve the entity-level information, Graph-based RAG approaches (Edge et al., 2024; Guo et al., 2025; Jimenez Gutierrez et al., 2024; Peng et al., 2024; Zhang et al., 2025; Han et al., 2024; gutiérrez2025ragmemory), such as GraphRAG (Edge et al., 2024) and LightRAG (Guo et al., 2025), propose to index the unstructured texts into the structured graphs. These approaches extract entities and relations from texts as graph nodes and edges, and their retrievals are based on the graph components mapped from the query text. However, existing Graph-based RAG approaches (Edge et al., 2024; Guo et al., 2025; Jimenez Gutierrez et al., 2024) still have three issues in the DFSTC task.
  1. (1)
    Uniform-indexing issue: Some of these approaches construct graphs by indexing extracted text chunks uniformly. They do not consider the varying importance of each text and extraction confidence of keyword entity. For example, in Figure 1, keyword lazy is related to all round 2 labels with different importance. Hence, uniform-indexing may provide inaccurate graph context for LLM classification.
  2. (2)
    Non-adaptive Retrieval issue: These approaches select relevant information for each input text based on a globally predefined threshold. However, the optimal retrieval threshold can vary across different query text and different rounds; hence, this non-adaptive retrieval approach can be suboptimal.
  3. (3)
    Narrow-source issue: These approaches only retrieve side information from a limited number of few-shot training text and ignore the important information among query texts. Since the new keywords related to certain labels in hate speech or misinformation detection emerge frequently in social media, this issue consequently limits models’ retrieval comprehensiveness.
To address these issues, we propose a novel Graph-based Online Retrieval Augmented Generation framework for dynamic few-shot social media text classification, called GORAG. In general, GORAG constructs and maintains an adaptive online weighted graph by extracting side information from all target texts, and tailoring the graph retrieval for each input.
Firstly, GORAG constructs a weighted graph with keywords extracted from labeled texts as keyword nodes, and text labels as label nodes, to help the model better capture graph topology (Yang et al., 2024b; Zhao et al., 2024; Lin et al., 2024). To model relationships, GORAG employs an edge-weighting mechanism to assign different weights to edges, representing keywords’ importance and relevance to the respective text’s label. Secondly, GORAG constructs a minimum-cost spanning tree based on query keywords and retrieves the label nodes within it as candidate labels. Since this MST is determined solely by the constructed graph and the keywords of the query text, GORAG achieves adaptive retrieval without human-defined thresholds, ensuring precision for each input. Finally, GORAG applies an online indexing mechanism to enrich the graph with keywords from query texts. By calculating weights that reflect the semantic relevance of keywords to labels, GORAG effectively reduces the potential noise introduced by inaccurate LLM classification during real-time updates, ensuring the graph remains a robust and comprehensive retrieval source
We summarize our novel contributions as follows.
  • We present an online RAG framework for DFSTC tasks, namely GORAG, which consists of three steps: Graph Construction, Candidate Label Retrieval, and Text Classification, aiming to address the Uniform-indexing, Non-adaptive Retrieval, and Narrow-source issues.
  • We develop a novel graph edge weighting mechanism based on the text keywords’ importance within the text corpus, which enables our approach to effectively model the relevance between keywords and labels, thereby helping the more accurate retrieval.
  • To achieve adaptive retrieval, we formulate the candidate label retrieval problem which is akin to the NP-hard Steiner Tree problem, and provide an efficient and effective solution based on the greedy algorithm.
  • Extensive experiments from the social media contextual abuse benchmark and the web news misinformation detection benchmark demonstrate the effectiveness and efficiency of GORAG.
2. Preliminary and Related Works
We first introduce the preliminaries of Dynamic Few-shot Social media Text Classification (DFSTC) in Section 2.1 and then discuss the related works for DFSTC task in Section 2.2. The important notations used in this paper are listed in Table 7 in Appendix A.
2.1. DFSTC Task
Text classification (Kowsari et al., 2019; Minaee et al., 2021; Li et al., 2022) is a key task in real-world web for good applications, including social media hate speech detection (Geissler et al., 2025; Hee et al., 2024; Chen et al., 2024b; Djuric et al., 2015; Subramanyam et al., 2024; Mandal et al., 2024; Cao et al., 2024; Ji et al., 2024; Guo et al., 2024), misinformation/disinformation detection (Luo et al., 2024; Nesterov et al., 2024; Liu et al., 2025; Toughrai et al., 2025; Cinus et al., 2025; Ashkinaze et al., 2024; Chen et al., 2022; Shang et al., 2024; Ye et al., 2024). It involves assigning predefined labels $y\in\mathcal{Y}$ to text $t=(w_{1},\cdots,w_{|t|})$ based on its words $w_{i}$. Traditional approaches require fine-tuning with large labeled datasets (Devlin, 2018; Liu, 2019), which may not always be available. Recently, few-shot learning addresses this limitation by enabling models to classify social media texts into various classes with only a small number of labeled examples per class (Meng et al., 2018, 2020). On the other hand, dynamic classification (Parisi et al., 2019; Hanmo et al., 2024; Wang et al., 2024a) introduces an additional challenge where new classes are introduced over multiple rounds, requiring the model to adapt to new classes while retaining knowledge of previously seen ones, as frequently involve extensive human labeling for these new classes can be difficult. For example, in Figure 1, new and more fine-grained classes of hate speech, such as Identity Insults, can serve as complements to categories that already exists.
Combining these aspects, Dynamic Few-Shot Social media Text Classification (DFSTC) task allows the model to handle evolving classification tasks with minimal labeled data, and is better suited for real world scenarios (Xia et al., 2021). In DFSTC task, the model is provided with multiple rounds of new class updates. Specifically, in each round $r$, a set of new labels $\mathcal{Y}^{r}_{new}$ is introduced. The complete candidate label set for round $r$, denoted as $\mathcal{Y}^{r}$, is the union of all new labels from rounds $i=1$ to $r$: $\mathcal{Y}^{r}=\bigcup_{i=1}^{r}\mathcal{Y}^{i}_{new}$. In round $r$, the labeled training set for class labels in $\mathcal{Y}^{r}$ is denoted as $\mathcal{D}^{r}=\cup_{y_{i}\in\mathcal{Y}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$, where per class $y_{i}\in\mathcal{Y}^{r}$ only has $k$ labeled examples $\{t_{j},y_{i}\}_{j=1}^{k}$. Formally, at the $r$-th round, given the labels in $\mathcal{Y}^{r}$ and all labeled data $\mathcal{D}^{r}=\cup_{i=1}^{r}\mathcal{D}^{i}$, the target of DFSTC task is to learn a function $f^{r}_{\theta}$, which can learn scores for all target labels $\mathbf{s}^{r}={f}^{r}_{\theta}(t_{u},\mathcal{Y}^{r})\in[0,1]^{|\mathcal{Y}^{r}|}$ for the unseen text $t_{u}$. Then, we can get the predicted label $y^{r*}\in\mathcal{Y}^{r}$ for the unseen text $t_{u}$ as follows.
(1) $y^{r*}=\arg\max_{y^{r}\in\cup_{i=1}^{r}\mathcal{Y}^{i}}({f}_{\theta}^{r}(t_{u},\cup_{i=1}^{r}\mathcal{Y}^{i})[y^{r}])$
DFSTC is valuable for its ability to learn from limited labeled data, adapt to evolving class distributions, and address real-world scenarios where classes and data evolve over time.
2.2. DFSTC Models
Current Dynamic Few-Shot Social media Text Classification (DFSTC) models can be broadly categorized into NN-based models and RAG-based models, RAG-based models can be further categorized into Long Context RAG, and Graph-based RAG models.
2.2.1. NN-based models
NN-based models (Meng et al., 2018, 2020; Xia et al., 2021) generate additional data contrastively based on the few-shot labeled data and the text formed label names to train the NN-based classifier. However, due to the limited labeled data, the generated text data of these models can have very limited patterns, which makes them prone to overfitting (Liu et al., 2023, 2024a; Lei et al., 2023). In web for good applications such as hate speech detection, different kinds of hate speech contents are usually a minority compared with neutral contents, this highly unbalanced data distribution can further overfit the NN-based models and affect their performance (Liu et al., 2023).
Figure 2. An overview of GORAG in each round. In Part 1, GORAG constructs a weighted graph based on keywords extracted from the few-shot training data. In Part 2, GORAG performs adaptive graph retrieval and outputs the candidate labels, which is a subset of the original target label set. In Part 3, GORAG first classifies texts into one of the candidate labels, and then applies online graph indexing to update the graph with query text extracted entities that not yet existed in the graph.
2.2.2. Long Context RAG models
Recently, Large Language Model (LLM)-based models have undergone rapid development (Zhang et al., 2024b; Snyder et al., 2024; Zhong et al., 2024; Yan et al., 2024; Chen et al., 2024a; Ding et al., 2024). Notably, LLMs are inherently capable of inference without fine-tuning (Sun et al., 2021; Jiang et al., 2023; Touvron et al., 2023; Achiam et al., 2023; Yang et al., 2024a), making them originally suitable for few-shot tasks. However, the lack of fine-tuning can lead LLMs to generate unfaithful responses, as they lack task-specific knowledge, these incorrect answers are often referred to as hallucinations (Zhang et al., 2023). In web for good applications such as hate speech detection, hallucinations can lead to failure to detect harmful responses, making these harmful contents being spread through internet. To mitigate hallucinations, researchers try to retrieve text documents outside of LLM as side information to help the LLM inference, namely Long Context RAG models (Chen et al., 2023; Sarthi et al., 2024; Hu et al., 2024; Wu et al., 2024a; Cai et al., 2024; Oosterhuis et al., 2024; Yadav et al., 2024; Che et al., 2024; Pan et al., 2024; Jiang et al., 2024). Among them, prompt compressor models (Pan et al., 2024; Jiang et al., 2024), apply LLM’s generation perplexity on to filter out un-important tokens in the model input.
2.2.3. Graph-based RAG models
However, the retrieved contents of Long Context RAG models remain unstructured and lengthy, making the important fine-grained entity-level information be overwhelmed by noisy contexts (Edge et al., 2024; Guo et al., 2025), which still impedes the effectiveness or efficiency of LLM classification, leading to lost-in-the-middle issues (Liu et al., 2024b) for cheaper open-sourced LLMs, or high token cost for proprietary LLMs (Hurst et al., 2024). To address these issues, Graph-based RAG models were proposed (Peng et al., 2024; Zhang et al., 2024a; Wu et al., 2024b; Zhang et al., 2024c; Guo et al., 2025; Jimenez Gutierrez et al., 2024; gutiérrez2025ragmemory), these models convert text chunks from the side information document as graph nodes, and the relation between these text chunks as graph edges. These models then retrieve graph nodes and edges as the LLM classification context, better preserving more fine-grained, entity-level information. However, current Graph-based RAG models usually consider each graph edge uniformly and ignore the varying confidence and importance of relations between text chunks. Also, as the optimal retrieval threshold can vary across different data samples, their globally fixed threshold selected by humans may be suboptimal for the entire dataset.
3. Methodology
3.1. Framework Overview
To address the aforementioned Uniform-indexing, Non-adaptive Retrieval, and Narrow-source issues, we propose GORAG, a novel approach that achieves adaptive retrieval by extracting valuable side information from a minimum-cost spanning tree generated on the constructed graph. As shown in Figure 2, GORAG consists of three core parts, i.e., Offline Graph Indexing, Graph Retrieval, and Classification & Online Graph Indexing.
Part 1: Offline Graph Indexing. It targets to construct the weighted graph $\mathcal{G}^{r}_{new}(\mathcal{V}^{r}_{new},\mathcal{E}^{r}_{new},\mathcal{W}^{r}_{new})$ based on the labeled data $\mathcal{D}^{r}=\cup_{y_{i}\in\mathcal{Y}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$ at the $r$-th round. An example weighted graph created is shown in Figure 2, this graph will be used to provide retrieved-augmented information as contexts for query texts, enabling LLMs to better understand query texts and accurately classify them. Specifically, with the new labeled data $\mathcal{D}^{r}$ at the $r$-th round, we first extract keywords $\mathcal{K}^{r}\in\{t_{j}\}_{j=1}^{|\mathcal{D}^{r}|}$. Then, we will assign an edge $e^{r}_{v,y}\in\mathcal{E}^{r}_{new}$ between each keyword $v\in\mathcal{K}^{r}$ and their respective text’s label $y\in\mathcal{Y}^{r}$, where $\mathcal{Y}^{r}=\bigcup_{i=1}^{r}\mathcal{Y}^{i}_{new}$ denotes the union label set of all new labels $\mathcal{Y}^{i}_{new}$ from round $i\in\{1,2,\dots,r\}$. We then compute the weight $w^{r}_{v,y}\in\mathcal{W}^{r}_{new}$ for each edge $e^{r}_{v,y}\in\mathcal{E}^{r}_{new}$ based on the keyword’s importance within the text corpus and keyword relatedness to the label $y$. At last, we merge all graphs $\mathcal{G}^{r}_{new}(\mathcal{V}^{r}_{new},\mathcal{E}^{r}_{new},\mathcal{W}^{r}_{new})$ at each round $r$ as the full graph $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{W}^{r})$. More details can refer to Section 3.2.
Part 2: Graph Retrieval. The Graph Retrieval process maps the extracted keyword nodes $\mathcal{V}^{t}$ from the query text $t$ to the constructed weighted graph $\mathcal{G}^{r}$ and generates a minimum-cost spanning tree on $\mathcal{G}^{r}$ that includes all mapped keywords in $\mathcal{V}^{t}$. From this minimum-cost spanning tree (MST), the candidate label set $\hat{\mathcal{Y}}_{t}^{r}$ at the current round $r$ is obtained, which is a reduced subset of the original target label set $\mathcal{Y}^{r}$ at the current round $r$. As this MST solely depends on the graph itself and the text extracted keywords, we can eliminate any human-defined thresholds and achieve adaptive retrieval. For more details, please refer to Section 3.3.
Part 3: Classification & Online Graph Indexing. In the Classification part, we create LLM input with the query text $t$, the candidate labels $\hat{\mathcal{Y}}_{t}^{r}$ retrieved from the weighted graph $\mathcal{G}^{r}$, and descriptions $\mathcal{K}_{y_{i}}$ associated with each candidate label $y_{i}\in\hat{\mathcal{Y}}_{t}^{r}$ if available. The LLM will carry out classification with this input. After classification, the online graph indexing procedure dynamically adds keywords $\mathcal{V}_{notexist}^{t}$, which are extracted from the query text $t$ but are not in the existing graph $\mathcal{G}^{r}$, as new graph nodes to further enrich the graph. We then assign edge weights based on these newly added keywords’ importance within the text corpus and their relatedness to the predicted label $y^{*}_{t}$ for the query text $t$. By penalizing the edges between nodes and their semantically irrelevant labels, GORAG can reduce the error introduced by inaccurate classification. The Online Graph Indexing procedure improves the model classification ability by updating the graph in real-time with new keywords. For more details, please refer to Section 3.4. Due to space limitations, we put the time and space complexity analysis of GORAG to Appendix D.
3.2. Part 1: Offline Graph Indexing
In this subsection, we introduce the graph construction procedure of GORAG, which includes a novel edge-weighting mechanism to address the uniform indexing issue. The pseudo code of graph construction is shown in Algorithm 1 in Appendix C.
For each round $r$, given its respective labeled training texts $\mathcal{D}^{r}=\cup_{y_{i}\in\mathcal{Y}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$, GORAG first extracts text keywords $\mathcal{K}^{r}$ from $\mathcal{D}^{r}$. Specifically, it calculates the normalized TFIDF score (69) as the correlation score $CS(v,t)$ for texts’ each word and phrase with less than 3 words (Wang et al., 2025).
(2) $\vskip-4.015ptCS(v,t)=\frac{count(v,t)}{|t|log(|\mathcal{T}^{r}|+1)}\times log\frac{|\mathcal{T}^{r}|+1}{|t_{j}:v\in t_{j},\ t_{j}\in\mathcal{T}^{r}|+1},$
where $\mathcal{T}^{r}=\cup_{i=1}^{r}\mathcal{T}^{i}$ denotes all training and query texts seen so far, $count(v,t)$ is the number of times that the term $v$ appears in the text $t$, and $|t_{j}:v\in t_{j},\ t_{j}\in\mathcal{T}^{r}|$ denotes the number of texts in the corpus $\mathcal{T}^{r}$ that contain the keyword $v$. Here, we will extract words or phrases with $CS(v,t)>\tau$ as our keyword. We set $\tau=0.5$ following the keyword extraction setting in (Wang et al., 2025). These keywords serve as graph nodes, and we denote them as keyword nodes.
Also, GORAG incorporates new labels $\mathcal{Y}^{r}_{new}=\mathcal{Y}^{r}\setminus\mathcal{Y}^{r-1}$ at the $r$-th round as graph nodes, denoted as label nodes, where $\mathcal{Y}^{r}$ is the total label set at round $r$. Hence, the graph node set $\mathcal{V}^{r}_{new}$ at the $r$-th round can be obtained as follows.
(3) $\vskip-5.01874pt\mathcal{V}_{new}^{r}=\mathcal{K}^{r}\cup\mathcal{Y}^{r}_{new}.$
Then, GORAG adds edges between each keyword node $v\in\mathcal{K}^{r}$ to its corresponding label node $y\in\mathcal{Y}^{r}$, indicating that the keyword $v$ appears in texts associated with the label $y$.
Next, considering the keywords $\mathcal{K}^{r}$ from the labeled text are not uniformly related to the text’s label in $\mathcal{Y}^{r}$, we apply an edge weighting mechanism to assign a weight $w_{v,y}$ to each keyword-label edge between any two nodes $v$ and $y$. This weight can reflect the correlation between keywords and each label. Firstly, we apply the correlation score $CS(v,t)$ assigned by Equation (2) to measure the importance and relatedness of a particular keyword $v\in\mathcal{K}^{r}$ w.r.t. the text $t$ it extracted from.
As the keyword $v$ can be extracted from multiple texts with different labels in different rounds, the final edge weight $w_{v,y}^{r}$ of edge $e_{v,y}$ at the $r$-th round is calculated as an average of all weights from all seen texts with label $y$:
(4) $w_{v,y}^{r}=\frac{\sum_{t_{j}\in\mathcal{T}^{r,v,y}}1-CS(v,t_{j})}{|\mathcal{T}^{r,v,y}|},$
where $\mathcal{T}^{r,v,y}=\{t_{j}|v\in t_{j}\land t_{j}\in\mathcal{T}^{r,y}\}$ is the text that contains keyword $v$ and labeled $y$. We denote the generated weighted graph for $\mathcal{K}_{new}^{r}$ and $\mathcal{Y}_{new}^{r}$ at the $r$-th round as $\mathcal{G}^{r}_{new}(\mathcal{V}_{new}^{r},\mathcal{E}_{new}^{r},\mathcal{W}_{new}^{r})$, where $\mathcal{V}^{r}_{new}$, $\mathcal{E}^{r}_{new}$, and $\mathcal{W}^{r}_{new}$ denotes the node, edge and edge weight set respectively. Lastly, $\mathcal{G}^{r}_{new}$ will be merged into the graph from previous rounds $\mathcal{G}^{r-1}$ to form the full graph $\mathcal{G}^{r}$ at round $r$:
(5) $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{W}^{r}),$
$\mathcal{V}^{r}=\mathcal{V}^{r}_{new}\cup\mathcal{V}^{r-1},$ $\ \mathcal{E}^{r}=\mathcal{E}^{r}_{new}\cup\mathcal{E}^{r-1},\ \mathcal{W}^{r}=\mathcal{W}^{r}_{new}\cup\mathcal{W}^{r-1}.$
Particularly, we define $\mathcal{G}^{0}=\emptyset$ and $r\geq 1$. Also, to guarantee the graph connectivity of the resulting graph, we add edges between every new label node $y_{n}\in\mathcal{Y}_{new}^{r}$, and each old label node $y_{o}\in\mathcal{Y}^{r-1}$, the edge weight $w_{y_{n},y_{o}}^{r}$ of edge $e_{y_{n},y_{o}}$ at round $r$ is defined as the weighted average of all edge weights that link keyword nodes with label node $y_{n}$ or $y_{o}$:
(6) $\vskip-5.01874pt\mathcal{M}_{y}^{r}=\{v\ |\ v\in\mathcal{N}^{r}(y)\ \land\ v\notin\mathcal{Y}^{r}\}\vskip-5.01874pt$
(7) $\vskip-3.01125ptw_{y_{n},y_{o}}^{r}=\frac{1}{2}\Big(\frac{\sum_{v\in\mathcal{M}_{y_{n}}^{r}}w_{v,y_{n}}^{r}}{2\times|\mathcal{M}_{y_{n}}^{r}|}+\frac{\sum_{v\in\mathcal{M}_{y_{o}}^{r}}w_{v,y_{o}}^{r}}{2\times|\mathcal{M}_{y_{o}}^{r}|}\Big),$
where $\mathcal{N}^{r}(y)$ denote the neighbor nodes of label node $y$ in $\mathcal{G}^{r}$.
After merge, the graph $\mathcal{G}^{r}$ would be used for future retrieval, and be further updated by GORAG’s online indexing mechanism.
3.3. Part 2: Graph Retrieval
In this subsection, we introduce the graph retrieval procedure of GORAG. With the graph constructed in Part 1, GORAG can adaptively retrieve a set of candidate class labels with keywords extracted from query texts without any human-defined thresholds, addressing the threshold dependent issue.
To begin with, GORAG extracts keywords $\mathcal{V}_{test}^{t}$ for each query text $t\in\mathcal{T}_{test}$ in the same manner with Equation (2), then, $\mathcal{V}_{test}^{t}$ would be split into two subsets:
(8) $\mathcal{V}_{exist}^{t}\cup\mathcal{V}_{notexist}^{t}=\mathcal{V}_{test}^{t},\ \mathcal{V}_{exist}^{t}\cap\mathcal{V}_{notexist}^{t}=\emptyset,$
where $\mathcal{V}_{exist}^{t}$ and $\mathcal{V}_{notexist}^{t}$ denotes the keywords in $\mathcal{V}_{test}^{t}$ that already exist and not yet exist in $\mathcal{G}^{r}$ at current round $r$ respectively. Later, $\mathcal{V}_{exist}^{t}$ would be applied for achieving adaptive retrieval, and $\mathcal{V}_{notexist}^{t}$ would be applied for online indexing to further enrich the constructed graph (Further illustrated in Section 3.4).
To achieve the adaptive retrieval, GORAG tries to find the minimum cost spanning tree (MST) that contain all keyword nodes within $\mathcal{V}_{exist}^{t}$, and then retrieves all label nodes within the generated MST as candidate labels. The intuition behind this approach is that an MST spans the entire graph to cover all given nodes with the smallest possible spanning cost. Consequently, label nodes within the generated MST can be considered important for demonstrating the features of the given keyword node set. As the generation of an MST is a classical combinatorial optimization problem with an optimal solution determined solely by the set of given nodes. By generating the MST and retrieving all label nodes within it, we eliminate the need for any human-defined thresholds, achieving adaptive retrieval.
Table 1. Statistics of the tested datasets, where we divide their original labels into multiple rounds. We achieve a balanced testing set on the IFS-Rel datasets by assigning each label 40 testing samples.
Dataset Avg. Text Token R1 R2 R3 R4 Total
Testing data Label # Testing data Label # Testing data Label # Testing data Label # Testing data
CAD (Vidgen et al., 2021) 154 240 4 60 3 286 3 231 3 817
WOS (Kowsari et al., 2017) 200 2,417 32 2,842 53 2,251 30 1,886 18 9,396
IFS-Rel (Xia et al., 2021) 105 640 16 640 16 640 16 640 16 2,560
Definition 3.1 (Adaptive Candidate Label Generation Problem).
Given an undirected weighted connected graph $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{W}^{r})$, a set of keywords $\mathcal{V}_{exist}^{t}$ extracted from text $t$ that can be mapped to nodes in $\mathcal{V}^{r}$, and the target label set $\mathcal{Y}^{r}\in\mathcal{V}^{r}$ at the $r$-th round, our target is to find a set of labels nodes ${\mathcal{Y}}^{r}_{t}\in{\mathcal{Y}}^{r}$. Firstly, we identify a subgraph $\mathcal{G}^{r}_{t}(\mathcal{V}^{r}_{t},\mathcal{E}^{r}_{t},\mathcal{W}^{r}_{t})$ of $\mathcal{G}^{r}$ by minimizing the edge weight sum as follows.
$\min_{e^{r}_{u.v}\in\mathcal{E}^{r}_{t}}w^{r}_{u,v}$
(9) $s.t.\$$v\in\mathcal{V}^{r}_{t},\forall v\in\mathcal{V}_{exist}^{t}$
Then, since the subgraph node $\mathcal{V}^{r}_{t}$ set contains both keyword nodes and label nodes, we define the candidate label set $\hat{\mathcal{Y}}^{r}_{t}\in\mathcal{V}^{r}_{t}$ as the label nodes in the generated sub-graph for the text $t$ at round $r$.
Due to space limitations, we put the proof of the Adaptive Candidate Label Generation Problem to be NP-hard to Appendix B.
As the problem is NP-hard, it is infeasible to obtain the optimal result in polynomial time. Therefore, to solve this problem, we propose a greedy algorithm based on the Melhorn’s algorithm (Mehlhorn, 1988). Our algorithm generates the candidate label set $\hat{\mathcal{Y}}^{r}_{t}\subseteq\mathcal{Y}^{r}$ for text $t$ at the $r$-th round. The pseudo code of our algorithm is shown in Algorithm 2 in Appendix C , following the Melhorn’s algorithm, our algorithm can also achieve 2-Approximate.
Firstly, we calculate the shortest path between each keyword node $v$ in $\mathcal{V}_{exist}^{t}$ to all other nodes in $\mathcal{G}^{r}$ by calculating the minimum spanning tree $MST$ of $\mathcal{G}^{r}$ Here, the keyword nodes $v$ in $\mathcal{V}_{exist}^{t}$ serve as terminal nodes that determines the final generated candidate labels w.r.t. to the weighted graph $\mathcal{G}^{r}$. Secondly, we create a new auxiliary graph $H$ where the edges represent the shortest paths between the closest terminal nodes. Thirdly, we construct the minimum spanning tree $MST^{\prime}$ of the auxiliary graph $H$ , and then add the shortest paths between each two nodes in $MST^{\prime}$ to the Steiner Tree $ST$. Lastly our candidate labels $\hat{\mathcal{Y}}^{r}_{t}$ for text $t$ are calculated as the intersection of $ST$ and all target labels $\mathcal{Y}^{r}$ at round $r$: $ST\cap\mathcal{Y}^{r}$.
3.4. Part 3: Classification & Online Indexing
In this subsection, we first introduce GORAG’s classification and online graph indexing procedure, where it classifies the texts and further enriches the weighted graph by adding query text information to the created graph, addressing the narrow source issue. To begin with, we introduce how GORAG performs text classification based the LLM. Specifically, given each unlabeled query text $t$, GORAG utilizes LLM to predict its class label $y^{*}$ by constructing an input prompt $c_{t}$ as follows:
(10) $c_{t}=Concat(t,\mathcal{K}^{t},\hat{\mathcal{Y}}^{r}_{t},\mathcal{K}_{\hat{\mathcal{Y}}^{r}_{t}}).$
$c_{t}$ can be considered as the text concatenation of the extracted keywords $\mathcal{K}^{t}$ from the text $t$, and the candidate labels $\hat{\mathcal{Y}}^{r}_{t}$ obtained by Algorithm 2 in Appendix C, we omit the more complex input structures as they rarely improve performance, but lead to higher costs (Zhou et al., 2023; Liu et al., 2024b). Also, the $\mathcal{K}_{\hat{\mathcal{Y}}^{r}_{t}}=\{\mathcal{K}_{y_{i}}\}_{i=1}^{|\hat{\mathcal{Y}}^{r}_{t}|}$ and $\mathcal{K}_{y_{i}}$ are the representative keywords of each label $y_{i}\in\hat{\mathcal{Y}}^{r}_{t}$.
(11) $\mathcal{K}_{y_{i}}=LLM(p_{gen},\mathcal{D}^{r}),$
which were generated by LLM with the label’s semantic label name if available. Next, with a classification instruction prompt $p_{classify}$.
(12) $y^{*}_{t}=LLM(p_{classify},c_{t}),$
the LLM would try to select the best-suited label $y^{*}_{t}\in\mathcal{Y}^{r}$ to annotate the text $t$.
To fully leverage the text-extracted keywords, GORAG utilizes an online indexing mechanism to incrementally update keywords that do not yet exist in the weighted graph $\mathcal{G}^{r}$ at the $r$-th round to $\mathcal{G}^{r}$ based on the text $t$’s predicted label $y^{*}_{t}$. To be specific, each keyword node $v\in\mathcal{V}_{notexist}^{t}$ would be added to the original graph’s node set $\mathcal{V}^{r}$ and be assigned with an edge $e_{v,y}$ connecting it with the text $t$’s predicted label $y_{t}^{*}$:
(13) $\vskip-3.01125pt\mathcal{V}^{r}=\mathcal{V}^{r}\cup\mathcal{V}_{notexist}^{t},\ \mathcal{E}^{r}=\mathcal{E}^{r}\cup\mathcal{E}_{oi}^{t},$
where $\mathcal{E}_{oi}^{t}=\{e_{v,y_{t}^{*}}|v\in\mathcal{V}_{notexist}^{t}\}$ denotes the set of all newly assigned edges $e_{v,y}$ between keyword node $v$ and its predicted label $y_{t}^{*}$, for these newly assigned edges, their weight is calculated with the edge weighting mechanism illustrated in Equation (4). The edge weights can reflect the semantic relevance of keywords to labels, and reduce the noise introduced by inaccurate LLM classification.
4. Experiments
Table 2. Classification accuracy on CAD, WOS, and IFS-Rel benchmarks.
Dataset Category Model Round 1 Round 2 Round 3 Round 4
1-shot 5-shot 10-shot 1-shot 5-shot 10-shot 1-shot 5-shot 10-shot 1-shot 5-shot 10-shot
CAD NN-based Entailment (Xia et al., 2021) 0.2125 0.2000 0.1208 0.1700 0.1060 0.1634 0.1177 0.0836 0.0785 0.0857 0.0700 0.0529
Long Context RAG NaiveRAG (Gao et al., 2023) 0.1875 0.1375 0.0958 0.1500 0.1100 0.0766 0.0802 0.0563 0.0392 0.0600 0.0405 0.0281
LongLLMLingua (Jiang et al., 2024) 0.1667 0.2417 0.2417 0.1709 0.2167 0.2200 0.1370 0.1587 0.1552 0.1092 0.1359 0.1285
Graph-based RAG GraphRAG (Edge et al., 2024) 0.3958 0.3625 0.3292 0.3167 0.2933 0.2833 0.1587 0.1621 0.1706 0.1200 0.1236 0.1322
LightRAG (Guo et al., 2025) 0.3875 0.3583 0.3373 0.3133 0.2867 0.2867 0.1638 0.1502 0.1571 0.1175 0.1102 0.1188
HippoRAG2 (gutiérrez2025ragmemory) 0.3667 0.2250 0.1792 0.2953 0.2000 0.1458 0.1820 0.1032 0.0746 0.1435 0.0668 0.0535
GORAG 0.6125 0.6125 0.6125 0.5867 0.5833 0.5882 0.3993 0.4010 0.4001 0.3133 0.3146 03139
WOS NN-based Entailment (Xia et al., 2021) 0.3695 0.3823 0.4187 0.3994 0.4471 0.4222 0.4510 0.4857 0.4787 0.4030 0.4387 0.4442
Long Context RAG NaiveRAG (Gao et al., 2023) 0.3885 0.3904 0.3897 0.2267 0.2154 0.2187 0.1821 0.1475 0.1799 0.1653 0.1556 0.1649
LongLLMLingua (Jiang et al., 2024) 0.3806 0.3823 0.3901 0.2155 0.2202 0.2198 0.1770 0.1567 0.1608 0.1468 0.1382 0.1493
Graph-based RAG GraphRAG (Edge et al., 2024) 0.3852 0.3897 0.3906 0.2213 0.2197 0.2219 0.1816 0.1770 0.1786 0.1641 0.1634 0.1625
LightRAG (Guo et al., 2025) 0.3930 0.3806 0.3815 0.2202 0.2216 0.2145 0.1743 0.1767 0.1799 0.1625 0.1626 0.1632
HippoRAG2 (gutiérrez2025ragmemory) 0.4133 0.4245 0.4208 0.2552 0.2656 0.2666 0.2205 0.2210 0.2242 0.1906 0.1860 0.1915
GORAG 0.4866 0.4990 0.5134 0.4790 0.5109 0.5284 0.4736 0.4996 0.5234 0.4420 0.4717 0.4929
IFS-Rel NN-based Entailment (Xia et al., 2021) 0.3391 0.6046 0.5859 0.4008 0.5516 0.5828 0.3061 0.4703 0.5193 0.2971 0.4395 0.4895
Long Context RAG NaiveRAG (Gao et al., 2023) 0.7343 0.7344 0.7406 0.5394 0.5406 0.5469 0.4719 0.4719 0.4688 0.5233 0.4938 0.5000
LongLLMLingua (Jiang et al., 2024) 0.8750 0.8906 0.9219 0.5734 0.5922 0.5806 0.4703 0.4741 0.5225 0.5819 0.4903 0.5038
Graph-based RAG GraphRAG (Edge et al., 2024) 0.6406 0.7813 0.8125 0.4406 0.4688 0.4719 0.3828 0.3734 0.3141 0.4297 0.4422 0.3922
LightRAG (Guo et al., 2025) 0.7578 0.8281 0.8594 0.4906 0.5155 0.5156 0.3802 0.4010 0.4427 0.4531 0.4804 0.4610
HippoRAG2 (gutiérrez2025ragmemory) 0.7313 0.7375 0.7406 0.4875 0.5016 0.4938 0.4297 0.4328 0.4438 0.4984 0.5008 0.4953
GORAG 0.9688 0.9688 0.9688 0.5938 0.6172 0.6250 0.4844 0.4948 0.5344 0.6165 0.4988 0.5469
In this section, we evaluate GORAG’s performance on the DFSTC task. Section 4.1 introduces the experimental setup, while Section 4.2 presents results on effectiveness and efficiency. Section 4.3 covers ablation studies, and Section 4.4 discusses generalization. Due to space limitation, we put experiment details, such as hardware settings and case study in Appendix E.
4.1. Experiment Settings
4.1.1. Few-shot setting
In this paper, we employ 1-shot, 5-shot, and 10-shot settings for few-shot training, where each setting corresponds to using 1, 5, and 10 labeled training samples per class, respectively. We omit experiments with more than 10 labeled samples per class, as for WOS dataset, there are already over 1300 labeled training data under 10-shot setting.
4.1.2. Datasets
We select the CAD (Vidgen et al., 2021), WOS (Kowsari et al., 2017), and IFS-Rel (Xia et al., 2021) benchmark to test GORAG’s performance. The texts in CAD are social media posts that potentially have different kinds of hate speech and abuse. The texts in the IFS-Rel benchmark are segments from web news, and their classifications help detect misinformation. The texts from the WOS benchmark are abstracts or introductions from academic papers, as the WOS dataset contains a vast number of candidate labels and has a large word count for each text; these characteristics make it a challenging benchmark for text classification models. The statistics of these benchmarks are shown in Table 1.
Due to the space limitations, we mainly display our experinents on the 4 rounds split version of all datasets in Table 2, and further experiment results of 6 and 8 rounds are shown in Figure 4.
Ethical Disclaimer. This research involves sensitive content. All datasets are publicly available, and our study adheres to institutional ethical guidelines to ensure that no personally identifiable information is disclosed. Examples shown in this paper are for illustrative purposes only and do not reflect the authors’ views.
4.1.3. Evaluation Metrics
In this paper, aligning with previous DFSTC models (Xia et al., 2021; Meng et al., 2020), we use classification accuracy as the evaluation metric. Given that LLMs can generate arbitrary outputs that may not precisely match the provided labels, we consider a classification correct only if the LLM’s output exactly matches the ground-truth label name or label number.
4.1.4. Baselines
We compare six baselines from three types as follows. The details of selected baselines are in Appendix E.1.
  • One NN-based model: Entailment (Xia et al., 2021), which converts the task to binary classification, and creates entailment pairs with the labeled data for training.
  • Two Long Context RAG models: NaiveRAG (Gao et al., 2023) splits few-shot labeled data into chunks and retrieves the most query similar chunks to augment LLM classification. LongLLMLingua (Jiang et al., 2024) further compresses the split chunks of NaiveRAG by filtering out chunk tokens regarding their LLM generation perplexity.
  • Three Graph-based RAG models: GraphRAG (Edge et al., 2024) extracts entities and relations from few-shot training data, and uses LLM to summarize graph neighborhoods as communities for retrieval. LightRAG (Guo et al., 2025) drops the community-related procedure and solely retrieves the 1-hop neighborhood of the unlabeled text mapped entities on the graph. HippoRAG2 (gutiérrez2025ragmemory) retrieves the Top-k candidate labels and their training texts based on their Personalized PageRank (PPR) score on the graph.
Furthermore, as solely put the labeled text within the LLM context can greatly increase costs and affect LLM performance (Chen et al., 2023; Jiang et al., 2024; Liu et al., 2024b), we compare the performance of solely using LLM for classification with GORAG under 0-shot setting in Table 6.
4.1.5. Hyperparameter Settings
Following the exact setting in Entailment’s original paper (Xia et al., 2021), we train the Entailment model with the RoBERTa-large PLM for 5 epochs with a batch size of 16, a learning rate of $1\times 10^{-6}$, and the Adam optimizer (Kingma, 2014). For LongLLMLingua, we test the compression rate within 0.75,0.8,0.85{0.75,0.8,0.85} and select 0.8, as it achieves the best overall classification accuracy. For all RAG-based baselines, we set all their hyperparameter the same as in their original paper, if not further illustrated. For GraphRAG and LightRAG, we use their local search mode, as it achieves the highest classification accuracy on the most complex WOS dataset. For GraphRAG and NaiveRAG, we use the implementation from (20), which optimizes their original code and achieves better time efficiency while not affect the performance. For all other baselines, we use their official implementations.
4.2. Few-shot Experiments
(a) Time cost on WOS dataset.
(b) Time cost on CAD dataset.
Figure 3. Total time cost on 2 datasets under 1-shot setting.
4.2.1. Effectiveness Evaluation
In Table 2, we apply Qwen2.5-7B backbone for CAD and WOS benchmark. Due to cost and API risk control (failing to respond when detecting dirty words) reasons, we only use the GPT-4o backbone on the IFS-Rel news benchmark. As shown in Table 2, GORAG achieves the best classification accuracy over all dynamic rounds on all benchmarks. Based on the experiment results, we have the following observations:
  • Firstly, with the training data increase (R1 to R4), NN-based Entailment suffers from performance decrease on CAD datasets. This is because the CAD datasets have an unbalanced label distribution, as noted in prior studies (Liu et al., 2023, 2024a; Lei et al., 2023), NN-based models that apply data augmentation techniques are prone to overfit when the ground truth label distribution are unbalanced.
  • Secondly, NaiveRAG performs better than Entailment in round 1 of the WOS and IFS-Rel datasets by avoiding overfitting, but its long inputs hinder its ability to adapt to dynamic label updates, leading to a significant performance drop from R1 to R4.
  • Thirdly, Graph-based RAG baselines only slightly improve performance across datasets. They struggle with dynamic label updates due to issues like uniform indexing, threshold dependency, and narrow sources, leading to inaccurate retrievals.
  • Lastly, from R1 to R4, GORAG adapts better to dynamic label updates, addresses the issues of Graph-based RAG models, and achieve better performance than all tested baselines.
Table 3. Graph incremental trends and memory usage on IFS-Rel dataset round 1 with GPT-4o LLM.
Object 25% 50% 75% 100%
Nodes 1644 3238 (+1394) 4482 (+1244) 5299 (+817)
Edges 1771 3265 (+1494) 4597 (+1332) 5503 (+906)
Memory 13.3 KB 27.6 KB 37.0 KB 46.4 KB
4.2.2. Efficiency Evaluation
To evaluate the efficiency of GORAG, we compare the total time cost of GORAG and other baselines on the 4 round split version of WOS and CAD datasets with Qwen2.5-7B backbone, the results are shown in Figure 3(a) and 3(b). Although GORAG’s online indexing mechanism introduces a vast amount of knowledge into the weighted graph, we can still achieve comparable or better time efficiency with other baselines. We also list the token usage comparison between GORAG and other baselines in Table 4, and the graph’s memory usage and incremental trends in Table 3, where GORAG consumes fewer tokens than current RAG baselines, and only introduce less than 50KB memory cost compared with solely applying LLMs.
4.3. Ablation Study
To further analyze GORAG’s performance, we conduct an extensive ablation study under 1-shot setting with Qwen2.5-7B backbone, as RAG models show similar performance across shot settings. Specifically, we examine how different GORAG components impact classification performance using its variants. GORAG offline removes the online indexing mechanism. GORAG unit removes the edge weighting mechanism, every edge in this variant is assigned with weight 1. GORAG keyword only uses the keyword extracted from the text to create LLM classification input, rather than the whole text.
Table 4. Token cost on IFS-Rel dataset with 1-shot setting.
Model AVG Token per Query AVG Cost per Query (USD)
GPT-4o 634.6 \$0.005
LongLLMLingua 2771.3 \$0.021
NaiveRAG 4537.4 \$0.034
LightRAG 7866.3 \$0.059
GraphRAG 9296.3 \$0.070
GORAG 557.5 \$0.004
As shown in Table 5, firstly, GORAG offline achieves the worst accuracy, because it lacks a comprehensive retrieval source, making its generated candidate labels inaccurate, hence limits its performance. Secondly, GORAG unit outperforms GORAG offline, but still sub-optimal to GORAG, demonstrates the importance of modeling the varying importance between keywords and each label, which leads to accurate retrievals. Lastly, GORAG keywords demonstrates the important information for text classification are mostly text keywords, as it achieves comparable or better performance than other ablated baselines.
Table 5. Ablation experiments on WOS for GORAG’s variants.
Dataset Model Round
R1 R2 R3 R4
WOS GORAG offline&unit 0.2979 0.2287 0.2645 0.2249
GORAG offline 0.3063 0.2302 0.2455 0.2156
GORAG unit 0.4706 0.4394 0.4407 0.3899
GORAG keyword 0.4746 0.4606 0.4455 0.4030
GORAG 0.4862 0.4649 0.4814 0.4210
4.4. More Generalization Evaluations
4.4.1. Zero-shot Experiments
To further study GORAG’s 0-shot ability, we compare GORAG’s performance with the original powerful LLM GPT-4o (Hurst et al., 2024) without RAG approaches, and some RAG baselines that are applicable to the 0-shot scenario in Table 6. Due to cost constraints, we conducted experiments on the second largest IFS-Rel dataset by providing only the label names, without any labeled data to GORAG for creating the graph. In the 0-shot setting, we input the labels and their LLM generated descriptions as system prompt for each model, and GORAG would extract keywords from label descriptions, without using any information from labeled texts. Note that the Entailment model cannot be applied in this setting, as they require labeled data for training. For other RAG-based baselines, they are considered equivalent to their base LLMs in the 0-shot setting due to a lack of external information. According to Table 6, GORAG can still achieve better performance than all tested baselines.
4.4.2. Round Generalization Evaluation
To evaluate GORAG’s ability to generalize to a larger number of dynamic rounds, we conducted additional experiments on the 6 and 8 round split versions of the WOS dataset under 1-shot and 5-shot settings, as the WOS dataset contains more testing data and labels than CAD dataset, and more lengthy, complex texts than the IFS-Rel dataset. In Figure 4, the results for rounds 1 to 4 are obtained from the 4-round split version, rounds 5 and 6 are obtained from the 6-round split version, and rounds 7 and 8 are obtained from the 8-round split version of the WOS dataset. This setting can ensure a sufficient amount of testing data at each round, As illustrated in Figure 4, the classification accuracy of GraphRAG and NaiveRAG drops significantly from round 1 to 8, while GORAG maintains competitive classification accuracy as the number of rounds increases through 1 to 8, with both 1 and 5-shot setting, demonstrating the importance of a more precise retrieval.
Table 6. 0-shot experiments for GORAG and RAG baselines on IFS-Rel dataset under 0-shot setting.
Model R1 R2 R3 R4
GPT-4o (Hurst et al., 2024) 0.9219 0.5109 0.4063 0.5936
LongLLMLingua (Jiang et al., 2024) 0.9281 0.5297 0.4234 0.5806
LightRAG (Guo et al., 2025) 0.6718 0.3047 0.2135 0.0977
GraphRAG (Edge et al., 2024) 0.5625 0.1875 0.1458 0.0938
GORAG 0.9328 0.5406 0.4375 0.6063
(a) 1-shot setting.
(b) 5-shot setting.
Figure 4. Experiment result of WOS dataset under 1-shot and 5-shot settings for at most 8 rounds.
5. Conclusion
In this paper, we propose GORAG, a Graph-based Online Retrieval Augmented Generation framework for the Dynamic Few-shot Social media Text Classification (DFSTC) task. Extensive experiments on hate speech and misinformation detection benchmarks with different characteristics demonstrate the effectiveness of GORAG in classifying social media texts with only a limited number or even no labeled data. GORAG shows its effectiveness in adapting to the dynamic updates of target labels by adaptively retrieving candidate labels to filter the large target label set in each update round. For future work, we aim to further enhance GORAG’s performance and explore its application in more general scenarios.
6. Acknowledgements
Lei Chen’s work is partially supported by National Key Research and Development Program of China Grant No. 2023YFF0725100, National Science Foundation of China (NSFC) under Grant No. U22B2060, Guangdong-Hong Kong Technology Innovation Joint Funding Scheme Project No. 2024A0505040012, AOE Project AoE/E-603/18, Theme-based project TRS T41-603/20R, CRF Project C2004-21G, Key Areas Special Project of Guangdong Provincial Universities 2024ZDZX1006, Guangdong Province Science and Technology Plan Project 2023A0505030011, Guangzhou municipality big data intelligence key lab, 2023A03J0012, HKUST(GZ) - CMCC(Guangzhou Branch) Metaverse Joint Innovation Lab under Grant No. P00659, Hong Kong ITC ITF grant PRP/004/22FX, Zhujiang scholar program 2021JC02X170, HKUST-Webank joint research lab.
Appendix A Notations
Due to the space limitations, we put the summary table on important notations here in Table 7.
Table 7. Summary on Important Notations.
Notations Meanings
$t$ Text.
$w$ The edge weight.
$r$ The round of label updating.
$\tau$ The keyword extraction threshold.
$y\in\mathcal{Y}^{r}$ The target label at round $r$.
$k$ The number of labeled data provided per target label.
$f^{r}_{\theta}$ The function that learn all target labels’ score for texts.
$\mathbf{s}^{r}$ The score for all target labels at round $r$.
$y^{r*}$ The predicted label for the text at round $r$.
$\mathcal{G}^{r}(\mathcal{V},\mathcal{E})$ weighted graph with node set $\mathcal{V}$ and edge set $\mathcal{E}.$
$\mathcal{Y}^{r}$ New target labels at round $r$.
$\mathcal{D}^{r}$ The labeled training text for all new labels at round $r$.
$\mathcal{T}^{r}$ Unlabeled texts at round $r$.
$\mathcal{T}_{test}^{r}$ The query text for at round $r$.
$\mathcal{K}^{r}$ The extracted keyword set for new labels of round $r$.
$\mathcal{V}_{new}^{r}$ The set of new graph nodes to be added in round $r$.
$\mathcal{N}(\cdot)$ The neighbor set of a node in the weighted graph.
$p_{extract}$ The extraction instruction prompt.
$p_{gen}$ The generation instruction prompt.
$p_{classify}$ The classification instruction prompt.
$CS(\cdot)$ The correlation score between keywords and labels.
$w^{r}_{v,y}$ The weight of edge $e_{v,y}$ in round $r$.
$P_{u,v}$ The shortest path between node $u$ and node $v$.
$\mathcal{V}^{t}_{exist}$ The keywords extracted from $t$ and exist in graph.
$\mathcal{V}^{t}_{notexist}$ The keywords extracted from $t$ and not exist in graph.
$\mathcal{E}_{oi}^{t}$ Edges added based on text $t$.
Appendix B Proof
Here we prove our Adaptive Candidate Label Generation problem can be reduced from the NP-hard Steiner Tree problem (Su et al., 2020).
Theorem B.1.
The Adaptive Candidate Label Generation problem is NP-hard.
Proof.
To demonstrate that the Adaptive Candidate Label Generation problem is NP-hard, we provide a simple reduction of our problem from the Steiner Tree problem. Since the Steiner Tree problem is already proven to be NP-hard (Su et al., 2020), we show that there is a solution for the Steiner Tree problem if and only if there is a solution for our problem. Firstly, given a solution $S$ for our problem, we can construct a Steiner Tree by generating a minimum spanning tree for all nodes in $S$ on graph $\mathcal{G}^{r}$ then connecting all nodes from $\mathcal{V}_{exist}^{t}$ to their closest neighbor nodes in $S$. Secondly, any Steiner Tree $ST$ that contains any node $y\in\mathcal{V}^{r}_{t}$ is also a solution of our problem with $|\{y\in\mathcal{V}^{r}_{t}\}|$ nodes. Thus, we prove that the adaptive candidate type generation problem is NP-hard. ∎
Appendix C Pseudo Code of GORAG
In this section, we provide the pseudo code of GORAG’s graph construction and adaptive candidate type generation algorithm in 1 and 2, respectively.
Appendix D Complexity Analysis
Firstly, we analyze the time and space complexity of GORAG’s graph construction. For the text keyword extraction procedure, the time complexity would be $O(|\mathcal{D}^{r}|)$; If label names are available, the time complexity of generating label descriptions and extract label keywords would be $O(|\mathcal{Y}^{r}|)$; Calculating the TFIDF and indexing edges to graph would require $O(|\mathcal{V}_{new}^{r}||\mathcal{D}^{r}|)$ times, and merging graph requires $O(|\mathcal{Y}^{r}||\mathcal{Y}^{r-1}|)$. Hence, the total time complexity of GORAG’s graph construction at the $r$-th round would be:
$O(|\mathcal{Y}^{r}||\mathcal{Y}^{r-1}|+|\mathcal{V}_{new}^{r}||\mathcal{D}^{r}|)$.
[…]
The time complexity of GORAG’s adaptive candidate type generation algorithm is the same with the Mehlhorn algorithm, which is $O(|\mathcal{E}^{r}|+|\mathcal{V}^{r}|log|\mathcal{V}^{r}|)$ (MEHLHORN1988125); The time complexity of the online indexing mechanism would cost $|\mathcal{V}_{notexist}^{r}||\mathcal{T}_{test}^{r}|$; Hence, the total time complexity of GORAG’s adaptive retrieval can be denoted as
$O(|\mathcal{T}_{test}^{r}|(|\mathcal{E}^{r}|+|\mathcal{V}_{notexist}^{r}||\mathcal{T}_{test}^{r}|+|\mathcal{V}^{r}|log|\mathcal{V}^{r}|)).$
For the space complexity, $O(|\mathcal{E}^{r}|)$ to store the graph, and $O(|\mathcal{T}_{test}^{r}|)$ space is needed to store the testing corpus. Hence, the total space complexity of GORAG’s Retrieval and Classification procedure is
$O(|\mathcal{E}^{r}|+|\mathcal{T}_{test}^{r}|).$
Algorithm 1 Graph Construction Algorithm of GORAG at round $r$
[…]
1: Let $\mathcal{V}_{text}^{r}=\emptyset$, $\mathcal{V}_{label}^{r}=\emptyset$, $\mathcal{E}_{new}^{r}=\emptyset$
2:for each text $t\in\mathcal{T}^{r}$ do
3:  $\mathcal{V}_{text}^{r}=\mathcal{V}_{text}^{r}\cup KeywordExtract(t)\cup y_{t}$
4:end for
5:if label names are available then
6:  for each label $y_{i}\in\mathcal{Y}^{r}$ do
7:   Get some label related keywords $\mathcal{V}_{label}^{r}=LLM(p_{gen},\mathcal{D}^{r})\cup\mathcal{V}_{label}^{r}$.
8:  end for
9:end if
10: Let $\mathcal{V}_{new}^{r}=\mathcal{V}_{text}^{r}\cup\mathcal{Y}^{r}_{new}\cup\mathcal{V}_{label}^{r}$ be the new node set at the $r$-th round.
11:for each node $v\in\mathcal{V}_{new}^{r}$ do
12:  $\mathcal{E}_{new}^{r}=\mathcal{E}_{new}^{r}\cup e_{v,y_{i}}$
13:  $\mathcal{W}_{new}^{r}=\mathcal{W}_{new}^{r}\cup w_{v,y_{i}}^{r}$
14:end for
15: Let $\mathcal{G}^{r}_{new}=(\mathcal{V}_{new}^{r},\mathcal{E}_{new}^{r})$ be the newly constructed graph for at the $r$-th round.
16: Let $\mathcal{G}^{r}=Merge(\mathcal{G}^{r}_{new},\mathcal{G}^{r-1})$ be the final constructed graph at the $r$-th round.
17:return $\mathcal{G}^{r}$
Appendix E Experiment Details
E.1. Introduction of chosen baselines
As shown in Table 2, we compare GORAG’s performance with 7 baselines from 3 technical categories for few-shot experiments, they are NN-based Entailment (Xia et al., 2021), Long Context-based NaiveRAG (Gao et al., 2023), LongLLMLingua (Jiang et al., 2024), and Graph-based GraphRAG (Edge et al., 2024), LightRAG (Guo et al., 2025), and HippoRAG2 (gutiérrez2025ragmemory).
NN-based Models
  • Entailment (Xia et al., 2021): Entailment concatenates the text data with each of the label names to form multiple entailment pairs with one text sample, hence increasing the number of training data and enhance its finetuning of a RoBERTa PLM (Liu, 2019).
Long Context RAG Models
  • NaiveRAG (Gao et al., 2023): It acts as a foundational baseline of current RAG models. When indexing, it stores text segments in a vector database using text embeddings. When querying, retrieves the side information based on their vector similarity to the query.
  • LongLLMLingua (Jiang et al., 2024): LongLLMLingua is a instruction aware prompt compressor model, it applies LLM’s generation perplexity to filter out un-important tokens of the model input based on the retrieved side information and the task instruction.
Graph-based RAG Models
  • GraphRAG (Edge et al., 2024): It employs LLM for graph construction. When retrieval, it aggregates nodes into communities w.r.t. the query, and generates community reports to encapsulate global information from texts.
  • LightRAG (Guo et al., 2025): It calculates the vector similarity between query extracted entities and graph nodes, achieving a one-to-one mapping from entities to graph nodes, then retrieve these nodes’ graph neighborhood.
  • HippoRAG2 (gutiérrez2025ragmemory): It employs LLM for graph construction, and also add the text chunks as nodes in the graph. As a result, it directly retrieves Top-k text chunks based on their respective nodes’ PPR score.
Algorithm 2 Adaptive Candidate Type Generation Algorithm
[…]
1: Compute a minimum spanning tree $MST$ of the graph $\mathcal{G}$.
2: Let $\mathcal{V}_{term}=\mathcal{V}_{exist}^{t}\cap\mathcal{V}(MST)$ be the terminal nodes in MSTMST
3: Construct a weighted auxiliary graph $H$: $\mathcal{V}(H)=\mathcal{V}_{term}$
4:for each pair of terminals $u,v\in\mathcal{V}_{term}$ do
5:  Find the shortest path $P_{uv}$ in MSTMST from $u$ to $v$
6:  Let $w^{H}_{u,v}=\sum_{e\in P_{uv}}w_{e}$
7:end for
8: Compute a minimum spanning tree $MST^{\prime}$ of the auxiliary graph $H$
9: Let $ST=\emptyset$
10:for each edge $e_{u,v}\in MST^{\prime}$ do
11:  Add the shortest path $P_{uv}$ in $MST$ to STST
12:end for
13: Let $\hat{\mathcal{Y}}^{r}_{t}=ST\cap\mathcal{Y}^{r}$
14:return $\hat{\mathcal{Y}}^{r}_{t}$
E.2. Hardware Settings
All experiments are conducted on an Intel(R) Xeon(R) Gold 5220R @ 2.20GHz CPU and a single NVIDIA A100-SXM4-40GB GPU.
E.3. Case Study
In this section, we dig into a test case from the CAD social media abuse detection dataset to demonstrate the ability of GORAG to provide graph-grounded clues, to explain its classification result.
As shown in Figure 5, the keyword entity feminists and sexual predators are extracted from the input unlabeled text and mapped to the graph. GORAG then generates a Minimum Cost Steiner Tree through these mapped entities, which also spans through the candidate label Insult. The structure and involved nodes of this generated Steiner Tree (edges in red) serve as an explicit clue, helping the LLM to classify the text correctly as Insulting Content.
Figure 5. A case from the CAD dataset, where the generated Steiner Tree serves as an explicit clue, helping the LLM to classify the text correctly.